# Discrete minimizers of the interaction energy in collective behavior: a brief numerical and analytic review

### v1

@@ Abstract
We consider minimizers of the $N$-particle interaction potential energy and briefly review numerical methods used to calculate them. We consider simple pair potentials which are attractive at short distances and repulsive at long distances, focusing on examples which are sums of two powers. The range of powers we look at includes the well-known case of the Lennard-Jones potential, but we are also interested in less singular potentials which are relevant in collective behavior models. We report on results using the software GMIN developed by Wales and collaborators for problems in chemistry. For all cases, this algorithm gives good candidates for the minimizers for relatively low values of the particle number $N$. This is well-known for potentials similar to Lennard-Jones, but not for the range which is of interest in collective behavior. Standard minimization procedures have been used in the literature in this range, but they are likely to yield stationary states which are not minimizers. We illustrate numerically some properties of the minimizers in 2D, such as lattice structure, Wulff shapes, and the continuous large-$N$ limit for locally integrable (that is, less singular) potentials.
@@ 1 Introduction
Finding configurations with minimum potential energy for a set of particles is a common mathematical problem found across several fields. In its simplest version we consider a potential $V\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ and a number $N\geq 2$ of particles. The problem then consists in finding the configurations $X=\{x_{1},\dots,x_{N}\}\subseteq\mathbb{R}^{d}$ such that the energy
@@
is as small as possible. These configurations are called minimizers or ground states. In a broader sense, the problem also involves understanding the structure of these minimizers: their shape, the distribution of their points, and their asymptotic properties for large $N$. The energy (1.1) is clearly invariant by a translation of the points $\{x_{1},\dots,x_{N}\}$, and consequently any translation of a minimizer is also a minimizer. One usually removes this invariance by looking for minimizers with some fixed property, most often by fixing the mean position $\sum_{i=1}^{N}x_{i}$ to be 00. If $V$ is radially symmetric, then $E_{N}$ is also invariant by rotations of the points $\{x_{1},\dots,x_{N}\}$, and by their inversion $\{-x_{1},\dots,-x_{N}\}$ (and hence by any rigid motion of these points). In all cases, minimizers are the same if $V$ is substituted by $V+C$ for some constant $C$, so some normalization on $V$ is usually assumed (often, either $\lim_{|x|\to+\infty}V(x)=0$ or $V(0)=0$, but not necessarily).
@@ 2 Numerical methods
In general the energy $E_{N}$ from (1.1), as a function $E_{N}\colon\mathbb{R}^{Nd}\to\mathbb{R}\cup\{+\infty\}$, is not convex, and the global optimization problem of finding its minima cannot use the comparatively simpler methods of convex optimization. As mentioned before, there’s the problem of invariance by translations, easily avoided by fixing the mean position of particles. The problem of invariance by rotations is harder to avoid, and hence algorithms usually find one of any possible rotations of the minimizer. It can then be hard to decide whether two runs of the same algorithm have given a rotation of the same minimizer, or a fundamentally different minimizer. However, the deeper problem is that there are usually a huge number of stationary states for the energy $E_{N}$; that is, points $\{x_{1},\dots,x_{N}\}$ at which
$\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\nabla V(x_{i}-x_{j})=0\qquad\text{for all $j=1,\dots,N$.}$ (2.1)
Many of these are local minima, and runs of standard local minimization algorithms will most often fall into one of these states, and not get to a minimizer. Hence simple descent algorithms such as variants of conjugate gradient descent, or Newton-type methods like LBFGS, will almost surely not find minimizers even for small $N$. Global minimization algorithms must somehow avoid these “traps” by more clever search techniques. We will give a short overview of numerical methods used for this kind of problems, and then describe the algorithm we use, implemented in the program GMIN by D. Wales (Wales, ).
@@
As discussed, these biased methods have been tried mainly for potentials of Lennard-Jones type, motivated by problems in chemistry and crystallography. Since our aim is to use a good numerical algorithm for a wider range of potentials whose minimizers may behave very differently, an adaptation of these methods does not seem straightforward, especially for the cases which have a continuous limit such as those discussed in Section 4, where particles do not arrange as large-scale crystals. Instead, we use a technique known as basin-hopping, first reported for the Lennard-Jones potential in Wales and Doye (1997), and before used by Li and Scheraga (1987) for a different problem in chemistry. This technique has been in particular quite successful for Lennard-Jones type potentials. It is an unbiased algorithm in the sense that it does not assume any particular structure of the minimizers, and it seems a more appropriate algorithm for the range of potentials we are interested in. It works as follows: first, the energy $E_{N}(X)$ is substituted by an energy $\widetilde{E}_{N}(X)$, obtained by performing a suitable descent algorithm starting from the configuration $X$ (the Polak-Ribiere variant of the conjugate gradient algorithm is mentioned in Wales and Doye (1997)). That is: if we denote by $M(X)$ the result of running a descent algorithm from the starting configuration $X$, then
@@
Hence the new energy landscape $\widetilde{E}_{N}$ is piecewise constant function, which is constant in a region where our chosen descent algorithm stops at the same local minimum $Y$, and whose value in this region is $E_{N}(Y)$. Now several algorithms are used:
@@
We have applied this algorithm for different potentials, otherwise with no modifications. Some of the results are explored in the next sections, together with a short review of the expected behavior of minimizers. We would like to emphasize that while obtaining local minimizers is straightforward in almost all cases, obtaining good candidates for a minimizer is a hard numerical problem. As an example, the hexagonal-shaped minimizers for very singular potentials such as Lennard-Jones (see Fig 1) require some kind of global search algorithm as described above. Or consider for example a minimizer for the locally integrable potential (1.2) with $a=2$ and $b=0.5$ (Figure 2). Asymptotically as $N\to+\infty$, this density is known to converge to the density of a minimizer of the continuous energy (1.5) (see Section 4), but the way this happens seems to show several interesting phenomena. One sees that locally particles still seem to be arranged in triangular patterns; but since the overall shape must be round (since the support of the minimizer is strongly suspected to be a ball), these local triangular arrangements must undergo some kind of dislocation patterns in order to fit the global shape. In Figure 2 one can clearly appreciate dislocation curves along which triangular patterns with different orientations meet. These patterns are quite sensitive to the minimization algorithm used, and we do not know whether the plot in Figure 2 actually represents a minimizer. But an understanding of these patterns is certainly an interesting problem, and even its numerical study requires powerful algorithms which can handle a large number of particles and yield good approximations to the minimizer. For particle numbers over a few thousand it is likely that this is still out of reach, even numerically.
Figure 2: Our best numerical candidate for a minimizer of the potential (1.2) with $a=2$, $b=0.5$ and $10^{4}$ particles. (We thank David Wales for providing this result)
@@ 3 Crystallization phenomena
The convergence here is the weak-*** convergence of measures (convergence when tested against continuous, compactly supported functions). Or even stronger: does this take place for all sequences of minimizers, up to rigid movements of these minimizers?
@@
when this sum converges absolutely. Also, given a subset $X=(x_{1},\dots,x_{N})\subseteq\mathbb{R}^{d}$ we define its energy per particle by
@@
In the context of crystallization phenomena one always assumes (3.1). As a consequence, a sequence of configurations with $N$ particles such that the interparticle distances all diverge to +$\infty$+ ∞ has energy as close to 00 as we wish. Hence for a minimizer $X_{N}$ it must happen that $E_{N}(X_{N})\leq 0$, so if Q3 holds one must have $e(L)\leq 0$. In any case Q3 in particular implies that for some $C>0$
$E_{N}(X_{N})\geq-CN$ (3.6)
for all $N\geq 2$. Potentials $V$ satisfying (3.6) are called H-stable, or sometimes just stable in the literature; see for example Ruelle (1969, Section 3.2). Often (3.6), though seemingly weaker, actually implies that $e_{N}(X_{N})$ does have a limit as $N\to+\infty$. Since the terminology comes from statistical mechanics and thermodynamics, interest is placed on potentials for which thermodynamical quantities (such as the energy-per-particle) have a finite limit, which justifies the word stable. Minimizers for potentials which are not stable may still have different scaling limits which are also interesting, though their study has different motivations and is more recent; see Section 4 below.
@@
$\frac{1}{N^{1/d}}X_{N}:=\Big{\{}\frac{x_{1}}{N^{1/d}},\dots,\frac{x_{N}}{N^{1/d}}\Big{\}},$
@@
We have focused on the above questions because we are interested in the behavior of $N$-particle minimizers. However, there are many related aspects of these problems which involve infinite distributions of particles. Among these we mention the following: ones:
@@
    Among all lattices in $\mathbb{R}^{d}$, which ones minimize the energy per particle (3.5)?
@@
    Given a countably infinite set of points in $\mathbb{R}^{d}$, it is possible to define their mean energy per particle in a reasonable way, so that it is consistent with the energy per particle of a lattice given in (3.3). Among all infinite particle distributions, which ones minimize the energy per particle?
We do not elaborate on different mathematical interpretations of the problem, namely: minimization problems with fixed density; modified models with a background energy; statistical mechanical or quantum mechanical models. We refer the reader to the reviews mentioned above for more on these matters (Radin, 1987; Blanc and Lewin, 2015).
@@
then minimizers cannot have any two particles at a distance closer than $1$ (otherwise their energy would be +$\infty$+ ∞), and there is a reward for spheres which are touching. Hence a minimizer of this potential is just a configuration of $N$ hard spheres in $d$-dimensional space with the maximum total number of contact points among them. In dimension $2$ this problem was solved by Heitmann and Radin (1980); see also Luca and Friesecke (2018) for a more recent proof. But the problem is far from understood in dimensions $3$ and higher, with open questions even for low values of $N$ (above $11$). A popular review of the problem can be found in Hayes (2012), and see Meng et al. (2010); Hoy et al. (2012) for some recent results. Attempts have been made to relate the structure of these minimizers to those for the Lennard-Jones potential (Trombach et al., 2018), but a precise link remains unclear. It is also tempting to make a connection of the sticky hard sphere problem with that of sphere packing, which is the problem of finding a distribution of identical hard spheres in $\mathbb{R}^{d}$ with maximum possible density. This link has often been remarked, but as far as we know there are no rigorous results backing this up. There have been some surprising recent results in sphere packing by Viazovska (2017); Cohn et al. (2017), which give specific lattices in dimensions $8$ and $24$ which can be proved to be the densest possible sphere distributions in these dimensions. The proof relies on the existence of specially symmetric lattices in these dimensions, which are natural candidates for this densest packing.
@@
In contrast, the problem Q1 is not so well understood even in dimension $1$. The best available result seems to be the one by Gardner and Radin (1979), which does show the strong version of Q1 for some potentials including the Lennard-Jones one. The conditions for this to hold are more restrictive, and it seems a hard problem to generalize them to a wide class of potentials.
@@
For the sticky hard sphere potential (3.8) in dimension $2$ results are already interesting. A first proof that minimizers must be placed at the sites of a triangular lattice was given by Heitmann and Radin (1980), giving an answer to Q1 (and hence Q2) and Q3. For a specific piecewise linear potential not too far from (3.8), a positive answer to Q3 was also given by Radin (1981). Then MO was answered by Au Yeung, Friesecke, and Schmidt (2011) for a family of short-range potentials with $V$ compactly supported and close enough to the sticky sphere potential (3.8): the convergence (3.7) takes place up to a subsequence, and the measure $\mu$ is $\frac{2}{\sqrt{3}}\mathbbm{1}_{E}$, where $\mathbbm{1}_{E}$ is the characteristic of a finite-perimeter set $E$ of area ${\sqrt{3}}/{2}$. In Theil (2006), Q3 is answered positively for a certain family of potentials which allow similar growth conditions as the Lennard-Jones one. In addition, it is proved that with periodic boundary conditions, the triangular lattice is a minimizer. Adding a three-body potential (and hence for an energy not strictly of the form (1.1)), and still with periodic boundary conditions, E and Li (2008) followed the ideas in Theil (2006) to show that minimizers must have particles placed at the sites of a hexagonal periodic arrangement (the vertices of a honeycomb, which are not a lattice in the sense we are using in this paper). Hence in this case, a version of Q2 holds, but not Q1.

### v2

@@ Abstract
We consider minimizers of the $N$-particle interaction potential energy and briefly review numerical methods used to calculate them. We consider simple pair potentials which are repulsive at short distances and attractive at long distances, focusing on examples which are sums of two powers. The range of powers we look at includes the well-known case of the Lennard-Jones potential, but we are also interested in less singular potentials which are relevant in collective behavior models. We report on results using the software GMIN developed by Wales and collaborators for problems in chemistry. For all cases, this algorithm gives good candidates for the minimizers for relatively low values of the particle number $N$. This is well-known for potentials similar to Lennard-Jones, but not for the range which is of interest in collective behavior. Standard minimization procedures have been used in the literature in this range, but they are likely to yield stationary states which are not minimizers. We illustrate numerically some properties of the minimizers in 2D, such as lattice structure, Wulff shapes, and the continuous large-$N$ limit for locally integrable (that is, less singular) potentials.
@@ 1 Introduction
Finding configurations with minimum potential energy for a set of particles is a common mathematical problem found across several fields. In its simplest version we consider a potential $V\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ and a number $N\geq 2$ of particles. The problem then consists in finding the configurations $X=(x_{1},\dots,x_{N})\in\mathbb{R}^{Nd}$ such that the energy
@@
is as small as possible. These configurations are called minimizers [^1] In some contexts they are called global minimizers, but in the present paper “minimizer” means “global minimizer” unless we explicitly mention local minimizers. or ground states. In a broader sense, the problem also involves understanding the structure of these minimizers: their shape, the distribution of their points, and their asymptotic properties for large $N$. The energy (1.1) is clearly invariant by a translation of the points $(x_{1},\dots,x_{N})$, and consequently any translation of a minimizer is also a minimizer. One usually removes this invariance by looking for minimizers with some fixed property, most often by fixing the mean position $\sum_{i=1}^{N}x_{i}$ to be 00. If $V$ is radially symmetric, then $E_{N}$ is also invariant by rotations of the points $(x_{1},\dots,x_{N}$, and by their inversion $(-x_{1},\dots,-x_{N})$ (and hence by any rigid motion of these points). In all cases, minimizers are the same if $V$ is substituted by $V+C$ for some constant $C$, so some normalization on $V$ is usually assumed (often, either $\lim_{|x|\to+\infty}V(x)=0$ or $V(0)=0$, but not necessarily).
@@ 2 Numerical methods
In general the energy $E_{N}$ from (1.1), as a function $E_{N}\colon\mathbb{R}^{Nd}\to\mathbb{R}\cup\{+\infty\}$, is not convex, and the global optimization problem of finding its minima cannot use the comparatively simpler methods of convex optimization. As mentioned before, there’s the problem of invariance by translations, easily avoided by fixing the mean position of particles. The problem of invariance by rotations is harder to avoid, and hence algorithms usually find one of any possible rotations of the minimizer. It can then be hard to decide whether two runs of the same algorithm have given a rotation of the same minimizer, or a fundamentally different minimizer. However, the deeper problem is that there are usually a huge number of stationary states for the energy $E_{N}$; that is, points $(x_{1},\dots,x_{N})$ at which
$\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\nabla V(x_{i}-x_{j})=0\qquad\text{for all $i=1,\dots,N$.}$ (2.1)
Many of these are local minima, and runs of standard local minimization algorithms will most often fall into one of these states, and not get to a minimizer. Hence simple descent algorithms such as variants of conjugate gradient descent, or Newton-type methods like LBFGS (Liu and Nocedal, 1989), will almost surely not find minimizers even for small $N$. Global minimization algorithms must somehow avoid these “traps” by more clever search techniques. We will give a short overview of numerical methods used for this kind of problems, and then describe the algorithm we use, implemented in the program GMIN by D. J. Wales (Wales, ).
@@
As discussed, these biased methods have been tried mainly for potentials of Lennard-Jones type, motivated by problems in chemistry and crystallography. Since our aim is to use a good numerical algorithm for a wider range of potentials whose minimizers may behave very differently, an adaptation of these methods does not seem straightforward, especially for the cases which have a continuous limit such as those discussed in Section 4, where particles do not arrange as large-scale crystals. Instead, we use a technique known as basin-hopping, first reported for the Lennard-Jones potential in Wales and Doye (1997), and before used by Li and Scheraga (1987) for a different problem in chemistry. This technique has been in particular quite successful for Lennard-Jones type potentials. It is an unbiased algorithm in the sense that it does not assume any particular structure of the minimizers, and it seems a more appropriate algorithm for the range of potentials we are interested in. It works as follows: first, the energy $E_{N}(X)$ is substituted by an energy $\widetilde{E}_{N}(X)$, obtained by performing a suitable descent algorithm starting from the configuration $X$ (the Polak-Ribiere variant of the conjugate gradient algorithm is mentioned in Wales and Doye (1997), although it currently uses a customized LBFGS method). That is: if we denote by $Y=M(X)$ the result of running a descent algorithm from the starting configuration $X$, then
@@
Hence the new energy landscape $\widetilde{E}_{N}$ is a piecewise constant function, which is constant in a region where our chosen descent algorithm stops at the same local minimum $Y$, and whose value in this region is $E_{N}(Y)$. This new landscape can be studied theoretically, taking $M(X)$ to be the point to which the descent algorithm converges, and it can be approximated numerically in practice by considering the point at which the numerical descent stops. Now several steps are followed:
@@
We have applied this algorithm for different potentials, otherwise with no modifications. Some of the results are explored in the next sections, together with a short review of the expected behavior of minimizers. We would like to emphasize that while obtaining local minimizers is straightforward in almost all cases, obtaining good candidates for a minimizer is a hard numerical problem. As an example, the hexagonal-shaped minimizers for very singular potentials such as Lennard-Jones (see Fig 1) require some kind of global search algorithm as described above. Or consider for example a minimizer for the locally integrable potential (1.2) with $a=2$ and $b=0.5$ (Figure 2). Asymptotically as $N\to+\infty$, this density is known to converge to the density of a minimizer of the continuous energy (1.5) (see Section 4), but the way this happens shows several interesting phenomena. One sees that locally particles still seem to be arranged in triangular patterns; but since the overall shape must be round (since the support of the minimizer is strongly suspected to be a ball), these local triangular arrangements must undergo some kind of dislocation patterns in order to fit the global shape. In Figure 2 one can clearly appreciate dislocation curves along which triangular patterns with different orientations meet. These patterns are quite sensitive to the minimization algorithm used, and we do not know whether the plot in Figure 2 actually represents a minimizer. But an understanding of these patterns is certainly an interesting problem, and even its numerical study requires powerful algorithms which can handle a large number of particles and yield good approximations to the minimizer. For particle numbers over a few thousand it is likely that this is still out of reach, even numerically. In Figures 2 and 3 we show minimizer candidates for 10,000 particles, which have required large computational power and several weeks of computer time.
Figure 2: Our best numerical candidate for a minimizer of the potential (1.2) with $a=2$, $b=0.5$ and $10^{4}$ particles. (We thank David J. Wales for providing this result)
@@ 3 Crystallization phenomena
The convergence here is the weak-* convergence of measures (convergence when tested against continuous, compactly supported functions). Or even stronger: does this take place for all sequences of minimizers, up to rigid movements of these minimizers?
@@
when this sum converges absolutely. Also, given a configuration $X=(x_{1},\dots,x_{N})\in\mathbb{R}^{Nd}$ we define its energy per particle by
@@
In the context of crystallization phenomena one always assumes (3.1). As a consequence, a sequence of configurations with $N$ particles such that the interparticle distances all diverge to +$\infty$+ ∞ has energy as close to 00 as we wish. Hence for a minimizer $X_{N}$ it must happen that $E_{N}(X_{N})\leq 0$, so if Q3 holds one must have $e(L)\leq 0$. In any case, if Q3 holds for the full sequence of minimizers (and not up to a subsequence), in particular it implies that for some $C>0$
$E_{N}(X_{N})\geq-CN,\quad\text{for all}\;N\geq 2.$ (3.6)
Potentials $V$ satisfying (3.6) are called H-stable, or sometimes just stable in the literature; see for example Ruelle (1969, Section 3.2). Often (3.6), though seemingly weaker, actually implies that $e_{N}(X_{N})$ does have a limit as $N\to+\infty$. Since the terminology comes from statistical mechanics and thermodynamics, interest is placed on potentials for which thermodynamical quantities (such as the energy-per-particle) have a finite limit, which justifies the word stable. Minimizers for potentials which are not stable may still have different scaling limits which are also interesting, though their study has different motivations and is more recent; see Section 4 below.
@@
$\frac{1}{N^{1/d}}X_{N}:=\left(\frac{x_{1}}{N^{1/d}},\dots,\frac{x_{N}}{N^{1/d}}\right),$
@@
We have focused on the above questions in order to narrow the discussion to a representative case, since the field is too broad for a short review. However, there are many related aspects of these problems. Some of them involve infinite distributions of particles. Among these we mention the following:
@@
    (Lattice minimization) Among all lattices in $\mathbb{R}^{d}$, which ones minimize the energy per particle (3.5)?
@@
    (Infinite particle distributions) Given a countably infinite set of points in $\mathbb{R}^{d}$, it is possible to define their mean energy per particle in a reasonable way, so that it is consistent with the energy per particle of a lattice given in (3.3). Among all infinite particle distributions, which ones minimize the energy per particle?
We do not elaborate on different mathematical interpretations of the problem, namely: minimization problems with fixed density; modified models with a background energy; statistical mechanical or quantum mechanical models. We refer the reader to the reviews mentioned above for more on these matters (Radin, 1987; Blanc and Lewin, 2015). For minimization on compact submanifolds of $\mathbb{R}^{d}$ we refer to Hardin and Saff (2005); Borodachov et al. (2019).
@@
then minimizers cannot have any two particles at a distance closer than $1$ (otherwise their energy would be +$\infty$+ ∞), and there is a reward for spheres which are touching. Hence a minimizer of this potential is just a configuration of $N$ hard spheres in $d$-dimensional space with the maximum total number of contact points among them. In dimension $2$ this problem was solved by Heitmann and Radin (1980); see also Luca and Friesecke (2018) for a more recent proof. But the problem is far from understood in dimensions $3$ and higher, with open questions even for low values of $N$ (above $11$). A popular review of the problem can be found in Hayes (2012), and see Meng et al. (2010); Hoy et al. (2012) for some recent results. Attempts have been made to relate the structure of these minimizers to those for the Lennard-Jones potential (Trombach et al., 2018), but a precise link remains unclear. It is also tempting to make a connection of the sticky hard sphere problem with that of sphere packing, which is the problem of finding a distribution of identical hard spheres in $\mathbb{R}^{d}$ with maximum possible density. This link has often been remarked, but as far as we know there are no rigorous results backing this up. There have been some surprising recent results in sphere packing by Viazovska (2017); Cohn et al. (2017), who give specific lattices in dimensions $8$ and $24$ which can be proved to be the densest possible sphere distributions in these dimensions. The proof relies on the existence of specially symmetric lattices in these dimensions, which are natural candidates for this densest packing.
@@
In contrast, problem Q1 is not so well understood even in dimension $1$. The best available result seems to be the one by Gardner and Radin (1979), which does show the strong version of Q1 for some potentials including the Lennard-Jones one. The conditions for this to hold are more restrictive, and it seems a hard problem to generalize them to a wide class of potentials.
@@
For the sticky hard sphere potential (3.8) in dimension $2$ results are already interesting. A first proof that minimizers must be placed at the sites of a triangular lattice was given by Heitmann and Radin (1980), giving an answer to Q1 (and hence Q2) and Q3. For a specific piecewise linear potential not too far from (3.8), a positive answer to Q3 was also given by Radin (1981). Then MO was answered by Au Yeung, Friesecke, and Schmidt (2011) for a family of short-range potentials with $V$ compactly supported and close enough to the sticky sphere potential (3.8): the convergence (3.7) takes place up to a subsequence, and the measure $\mu$ is $\frac{2}{\sqrt{3}}\mathbbm{1}_{E}$, where $\mathbbm{1}_{E}$ is the characteristic of a finite-perimeter set $E$ of area ${\sqrt{3}}/{2}$. In Theil (2006), Q3 is answered positively for a certain family of potentials which allow similar growth conditions as the Lennard-Jones one. In addition, it is proved that with periodic boundary conditions, the triangular lattice is a minimizer. Recently in Bétermin (2023) a computer-assisted method is used to show that the triangular lattice minimizes energy among lattices for the Lennard-Jones potential in 2D (hence giving an answer to the related lattice minimization problem we mentioned earlier).
Adding a three-body potential (and hence for an energy not strictly of the form (1.1)), and still with periodic boundary conditions, E and Li (2008) followed the ideas in Theil (2006) to show that minimizers must have particles placed at the sites of a hexagonal periodic arrangement (the vertices of a honeycomb, which are not a lattice in the sense we are using in this paper). Hence in this case, a version of Q2 holds, but not Q1. Problem Q3 has been solved by Bétermin et al. (2021) for some specific piecewise-constant potentials which give rise to a square lattice.